# ROTATE: Regret-driven Open-ended Training for Ad Hoc Teamwork

### v1

@@ Abstract
Developing AI agents capable of collaborating with previously unseen partners is a fundamental generalization challenge in multi-agent learning, known as Ad Hoc Teamwork (AHT). Existing AHT approaches typically adopt a two-stage pipeline, where first, a fixed population of teammates is generated with the idea that they should be representative of the teammates that will be seen at deployment time, and second, an AHT agent is trained to collaborate well with agents in the population. To date, the research community has focused on designing separate algorithms for each stage. This separation has led to algorithms that generate teammate pools with limited coverage of possible behaviors, and that ignore whether the generated teammates are easy to learn from for the AHT agent. Furthermore, algorithms for training AHT agents typically treat the set of training teammates as static, thus attempting to generalize to previously unseen partner agents without assuming any control over the distribution of training teammates. In this paper, we present a unified framework for AHT by reformulating the problem as an open-ended learning process between an ad hoc agent and an adversarial teammate generator. We introduce ROTATE, a regret-driven, open-ended training algorithm that alternates between improving the AHT agent and generating teammates that probe its deficiencies. Extensive experiments across diverse AHT environments demonstrate that ROTATE significantly outperforms baselines at generalizing to an unseen set of evaluation teammates, thus establishing a new standard for robust and generalizable teamwork.
@@ 1 Introduction
As the deployment of AI agents in diverse applications becomes more common, it is increasingly crucial that they can collaborate effectively with previously unseen humans and other AI agents. While methods for training teams of agents have been explored in cooperative multi-agent reinforcement learning (CMARL) [18, 43], previous work [53, 40] highlights that CMARL-based agents fail to perform optimally when collaborating with unfamiliar teammates. Rather than learning strategies that are only effective against jointly trained teammates, dealing with previously unseen teammates requires adaptive AI agents that efficiently approximate the optimal strategy for collaborating with diverse teammates. Methods to train adaptive agents in the context of collaborative tasks have previously been explored within the literature of ad hoc teamwork (AHT) [6, 50, 34] and zero-shot coordination (ZSC) [21, 13, 33].
@@
Prior work [34] has often decomposed AHT learning into two stages, consisting of first creating a fixed set of teammates, and then training an AHT agent using reinforcement learning (RL) approaches, based on interactions with teammates sampled from the set. Despite relying on neural network policies, algorithms that train an AHT agent based on interaction with a human-designed set of heuristic or pretrained agents [38, 64, 40] often struggle to handle novel behaviors outside the predefined set of teammates [51, 8]. Recent work [33, 9, 41, 42, 46] strengthens the generalization capabilities of previous methods by substituting the human-designed set of teammates with a set of teammate policies that are trained to maximize different notions of diversity. One such diversity notion is adversarial diversity [41, 9], which seeks to generate a set of teams that cooperate well within the team, but not across teams. However, prior work [14, 46, 10] empirically demonstrates that adversarial diversity often leads to teammate policies that actively diminish returns when interacting with agents other than their identified teammate, a phenomenon sometimes called self-sabotage.
@@
Figure 1: ROTATE Overview. ROTATE is an open-ended ad hoc teamwork learning framework in which the AHT agent learns as the set of training teammates expands. The core idea of ROTATE is to improve the AHT agent and iteratively generate diverse teammates with whom the AHT agent struggles to collaborate, yet not so adversarial that effective teamwork becomes impossible.
@@
This paper addresses two issues that cause existing AHT and ZSC methods to fail to learn policies that effectively collaborate with some teammates. First, existing methods [38, 64, 40, 41] learn from sampling teammates from a fixed set containing few teammate policies. Compared to the vast space of strategies a teammate may adopt, the AHT agent will only be trained to collaborate optimally with a small set of strategies, while potentially performing poorly against others. Second, existing methods focus on designing diverse and incompatible teammates [9, 41, 62], whose return-diminishing tendencies make it challenging for a randomly initialized, RL-based AHT agent to effectively learn to collaborate with.
In this paper, we present a fresh perspective on AHT, by observing that maximizing the return of an AHT agent on an unknown set of teammates is equivalent to minimizing its cooperative regret: the utility gap between the best response to a given teammate, and the AHT agent’s performance with that teammate. Inspired by the success of regret for designing generally capable agents that efficiently solve a broad range of tasks [56, 15, 24, 44], we then reformulate the AHT problem as a minimax game between the AHT agent and a teammate generator. Our problem formulation suggests an open-ended framework for AHT, that drives a teammate generator to continually discover new teammate policies, while jointly improving the AHT agent. Building on this theoretical foundation, we propose a practical algorithm, ROTATE (Fig. 1), which optimizes a regret-based minimax objective for both players, while maintaining a population of all teammates explored. Key to the success of ROTATE is a novel and practical per-state regret objective, designed to mitigate the self-sabotage problem that naturally arises from cooperative regret type objectives. We demonstrate that ROTATE significantly improves the robustness of AHT agents when faced with previously unseen teammates, compared to a range of baselines on Level-Based Foraging and Overcooked tasks.
@@ 2 Related Work
Agent Training in AHT & ZSC. The training of adaptive ego agent policies that can near-optimally collaborate with diverse previously unseen teammates has been explored in AHT [50] & ZSC [21]. Given access to a set of training teammate policies, AHT methods [34] train ego agents to model teammates [3] by identifying important characteristics of different teammates based on their observed behavior. These methods then train a model estimating the best-response policy to the encountered teammate policies based on their inferred characteristics. Recent AHT methods [40, 38, 64, 20, 55] typically use neural networks trained using reinforcement learning [48, 35]. To further improve AHT training, several approaches learn a distribution for sampling teammate policies during training based on maximizing the worst-case returns of trained agents [54] or maximizing regret [17, 11]. As an alternative to AHT, ZSC designs learning methods promoting near-optimal collaboration between agents that have not interacted with each other as long as they learn using the same ZSC algorithm. ZSC methods [21, 22, 13] typically achieve this goal by encouraging the agents to converge towards the same equilibrium despite being trained independently. These agent training algorithms are a crucial component of ROTATE, an iterative training procedure that uses AHT training algorithms to improve an AHT policy based on an interaction with generated teammate policies.
Teammate Generation for AHT & ZSC. Recent work improves existing AHT and ZSC agent training algorithms by designing a diverse collection of training teammate policies. FCP [51] generates the pretrained teammate policies by running the same CMARL algorithm across different seeds. Other work improved FCP by optimizing information-theoretic diversity metrics based on Jensen-Shannon divergence [33], mutual information [32], or entropy [59, 63], which encourages each teammate to yield different trajectories or policies. Recent methods [41, 9, 14, 42, 46] enable the trained AHT agent to learn distinct strategies during training by generating teammate policies requiring distinct best response policies through the maximization of adversarial diversity metrics, which strongly resembles ROTATE’s notion of cooperative regret. However, instead of maximizing the regret of the trained AHT agent like ROTATE, these methods maximize the regret from using a generated teammate policy to interact with another generated teammate policy. Unlike ROTATE’s open-ended training process, these methods also only generate a limited and fixed set of teammate policies once before agent training. Notably, these methods maximize regret only on the initial state of a collaboration trajectory, leading to sabotaging teammate policies [14, 46] that execute detrimental actions for cooperation in states that it will not visit in self-play. Learning to collaborate with sabotaging teammates is difficult, leading to the proposal of heuristics to reduce sabotage in previous work [14, 46], and a more systematic objective in ROTATE.
Open-Ended Learning (OEL). Our proposed method is heavily influenced by prior work in OEL [28, 52], which explores algorithms that continually design novel tasks to create a generally capable agent [23, 4]. Many OEL methods in RL [56, 15, 39, 25, 24, 44] focus on the problem of unsupervised environment design (UED), which aims to improve RL agent generalization across different tasks by designing and sampling novel environments with different transition and reward functions. Similar to our method, PAIRED [15] trains a neural network using reinforcement learning to output novel environment parameters that induce high regret to the trained ego agent. Other methods assume access to a procedural environment generator, and focus on designing task curators that sample training environments based on criteria such as the expected returns of different policies [56], TD-Error induced during learning [25], regret [24], or learnability [44]. In the context of competitive multi-agent RL, OEL methods seek to generate new opponents for competitive gameplay, often through self-play [49, 29]. For AHT, Yuan et al. [62] and Yan et al. [60] also proposed open-ended methods that keep generating novel teammate policies for an AHT agent to learn from. Unlike ROTATE, their approach to generating teammates either relies on evolutionary methods to generate new teammates or uses random perturbations of the AHT agent’s policy as the new teammate, making it less efficient at producing representative samples from the vast teammate policy set.
@@ 3 Background
The interaction between agents in an AHT setting may be modeled as a decentralized, partially observable Markov decision process (Dec-POMDP) [37]. A Dec-POMDP is characterized by a 9-tuple, $\langle N,S,\{\mathcal{A}^{i}\}_{i=1}^{|N|},P,p_{0},R,\{\Omega^{i}\}_{i=1}^{|N|},O,\gamma\rangle$, where $N$, $S$, and $\gamma$ respectively denote the index set of agents within an interaction, the state space, and a discount rate in $(0,1)$. Each interaction between the AHT agent and its teammates begins from an initial state $s_{0}$ sampled from an initial state distribution, $p_{0}(s)$. Denoting the set of all probability distributions over a set $X$ as $\Delta(X)$ and the current timestep as $t$, a Dec-POMDP assumes that each agent may not perceive the current state, $s_{t}$. Each agent instead perceives an observation from its observation space, $o^{i}_{t}\in\Omega^{i}$, sampled from the observation function, ${O:S\mapsto{\Delta(\Omega^{1}\times\dots\times\Omega^{|N|})}}$. Each agent $i\in{N}$ then chooses an action at time $t$ from its action space, $a^{i}_{t}\in\mathcal{A}^{i}$, based on a policy, $\pi^{i}(H^{i}_{t})$, conditioned on its observation-action history, $H^{i}_{t}=\{o^{i}_{\leq{t}},a^{i}_{<t}\}$. The actions selected by each agent are then collectively executed as the joint action, $a_{t}=(a^{1}_{t},\dots,a_{t}^{|N|})$. Each agent receives a common scalar reward, $r_{t}$, based on the reward function, $R:S\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{|N|}\mapsto\mathbb{R}$. Finally, a new state $s_{t+1}$, is sampled according to the environment transition function, $P:S\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{|N|}\mapsto\Delta{(S)}$. In this paper, the notation $\pi^{\text{ego}}$ refers to a trained AHT agent policy, or ego agent, while $\pi^{-i}$ refers to the $N-1$ policies of the AHT agent’s teammates. Importantly, we assume that teammates choose their actions only based on the current state. At the same time, the AHT agent selects its actions based on its state-action history, which is necessary to allow the AHT agent to distinguish between different types of teammates effectively.
@@ 4 Ad Hoc Teamwork Problem Formulation
Ad Hoc Teamwork (AHT) methods aim to train an adaptive AHT policy that an ego agent can follow to achieve maximal return when collaborating with an unknown set of evaluation teammates. Using the Dec-POMDP formulation to model the interaction between agents, this section formalizes the objective of AHT. While the most general AHT setting considers a possibly varying number of ego agents and teammates within an interaction [55, 40], this formalization addresses the more straightforward case where there is only a single ego agent within a team.
Let $\pi^{-i}$ denote a joint teammate policy controlling the $N-1$ non-ego agents during collaboration. Denote the returns of an ego agent following $\pi^{\text{ego}}$ to collaborate with teammates controlled by $\pi^{-i}$, starting from state $s$, as:
$V(s|\pi^{-i},\pi^{\text{ego}})=\mathbb{E}_{\begin{subarray}{c}a^{\text{ego}}_{t}\sim\pi^{\text{ego}},\\ a^{-i}_{t}\sim\pi^{-i},P,\,O\end{subarray}}\Bigg{[}\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\Bigg{|}s_{0}=s\Bigg{]}.$ (1)
@@
[…]
In some cases, an AHT algorithm can directly estimate this optimal policy by using $\Pi^{\text{eval}}$ to train an ego agent policy that maximizes $V(\psi^{\text{eval}},\Pi^{\text{eval}},\pi^{\text{ego}})$ when $\Pi^{\text{eval}}$ is known. [^1] In the context of reinforcement-learning-based AHT algorithms, “known” means that an AHT algorithm has unlimited sampling access to the teammate policies. However, most AHT methods address the more challenging case where $\Pi^{\text{eval}}$ is unknown, which is the setting that this paper adopts as well. While our methods assume no knowledge of $\Pi^{\text{eval}}$ during training, we follow standard practice [38, 40, 64, 55] by manually designing a diverse $\Pi^{\text{eval}}$ for evaluation purposes, as we later describe in Section 7.
When $\Pi^{\text{eval}}$ is unknown, AHT algorithms [34] either assume access to a training set of teammate policies, $\Pi^{\text{train}}$, or generate such a set. An expert’s domain knowledge about the characteristics of $\Pi^{\text{eval}}$ may be leveraged to construct a $\Pi^{\text{train}}$ similar to $\Pi^{\text{eval}}$. Once the set of training teammates has been formed, current AHT algorithms use reinforcement learning to discover an ego agent policy based on interactions with joint policies sampled from $\Pi^{\text{train}}$. While the precise training objective varies with the AHT algorithm, a common objective is to maximize the expected return during interactions with joint policies sampled uniformly from $\Pi^{\text{train}}$:
@@
Naturally, even an optimal ego agent policy, $\pi^{*,\text{ego}}(\Pi^{\text{train}})$, may not be optimal with respect to $\Pi^{\text{eval}}$ and $\psi^{\text{eval}}$, due to the potential distribution shift caused by differences between the training and evaluation objectives.

### v2

@@ Abstract
Learning to collaborate with previously unseen partners is a fundamental generalization challenge in multi-agent learning, known as Ad Hoc Teamwork (AHT). Existing AHT approaches often adopt a two-stage pipeline, where first, a fixed population of teammates is generated with the idea that they should be representative of the teammates that will be seen at deployment time, and second, an AHT agent is trained to collaborate well with agents in the population. To date, the research community has focused on designing separate algorithms for each stage. This separation has led to algorithms that generate teammates with limited coverage of possible behaviors, and that ignore whether the generated teammates are easy to learn from for the AHT agent. Furthermore, algorithms for training AHT agents typically treat the set of training teammates as static, thus attempting to generalize to previously unseen partner agents without assuming any control over the set of training teammates. This paper presents a unified framework for AHT by reformulating the problem as an open-ended learning process between an AHT agent and an adversarial teammate generator. We introduce ROTATE, a regret-driven, open-ended training algorithm that alternates between improving the AHT agent and generating teammates that probe its deficiencies. Experiments across diverse two-player environments demonstrate that ROTATE significantly outperforms baselines at generalizing to an unseen set of evaluation teammates, thus establishing a new standard for robust and generalizable teamwork.
@@ 1 Introduction
As AI agents are deployed in diverse applications, it is increasingly crucial that they can collaborate effectively with previously unseen AI agents and humans. While methods for training teams of agents have been explored in cooperative multi-agent reinforcement learning (CMARL) [20, 43], prior work highlighted that CMARL agents fail to perform optimally when collaborating with unfamiliar teammates [53, 40]. Rather than learning strategies that are only effective against jointly trained teammates, dealing with previously unseen teammates requires adaptive AI agents that efficiently approximate the optimal strategy for collaborating with diverse teammates. The training of such adaptive agents has been explored within ad hoc teamwork (AHT) [7, 50, 34] and zero-shot coordination (ZSC) [22, 15, 33].
@@
Most work has decomposed AHT learning into two stages [34], consisting of first creating a fixed set of teammates, and then training an AHT agent using reinforcement learning (RL), based on interactions with teammates sampled from the set. Methods that focus on AHT agent learning typically rely on a human-designed heuristic-based or pretrained teammates [39, 64, 40] and therefore struggle to handle novel behaviors outside the predefined set of teammates [51, 9]. Recent work enhances the generalization capabilities of AHT agent learning methods by substituting the predefined set of teammates with a generated collection of diverse teammates [33, 42], which are trained to maximize different notions of diversity. One such diversity notion is adversarial diversity [41, 10], which seeks to generate a set of teams that cooperate well within teams, but not across teams. However, prior work [16, 46, 11] empirically demonstrates that adversarial diversity often leads to teammate policies that actively diminish returns when interacting with agents other than their identified teammate, a phenomenon sometimes called self-sabotage.
@@
Figure 1: ROTATE Overview. ROTATE is an open-ended learning framework for AHT. The core idea of ROTATE is to improve the AHT agent by iteratively generating diverse teammates with whom the AHT agent struggles to collaborate, yet not so adversarial that effective teamwork becomes impossible.
@@
This paper addresses two issues that cause current methods to fail to learn policies that effectively collaborate with some teammates. First, two-stage AHT methods [39, 64, 40, 41] learn from interacting with teammates from a small fixed training set. Even when the training set is diverse, the AHT agent remains incapable of collaborating effectively with some teammates sampled from the vast space of possible strategies, specifically those with significantly different behavior from the policies in the training set [53, 40]. Second, other work designs a diverse training set of teammate policies by maximizing adversarial diversity [10, 41], which yields self-sabotaging teammates whose return-diminishing tendencies make it challenging for a randomly initialized RL-based AHT agent to learn to collaborate effectively [16, 46]. Despite addressing the first issue, some methods remain susceptible to the second issue by optimizing adversarial diversity [62].
We address the shortcomings of using a small, fixed training set by proposing an open-ended learning framework that continually generates new teammates with whom the AHT agent interacts to enhance its collaborative capabilities. We formulate our learning objective by observing that maximizing the expected returns of an AHT agent on a known set of teammates is equivalent to minimizing its expected cooperative regret: the utility gap between the best response to a given teammate, and the AHT agent’s performance with that teammate. While not knowing the teammates that will be encountered, we take inspiration from unsupervised environment design (UED) methods [56, 17, 24, 44] and train an AHT agent to minimize its regret against generated teammates that maximize the AHT agent’s cooperative regret. We propose a novel and practical objective that, unlike UED methods that optimize regret only at the initial state, also maximizes regret in states encountered later in an interaction. We build on these foundations to propose a practical algorithm, ROTATE (Fig. 1), which optimizes a cooperative regret-based minimax objective while maintaining a population of all teammates explored. We demonstrate that ROTATE significantly improves the robustness of AHT agents when faced with previously unseen teammates, compared to a range of baselines on two-player Level-Based Foraging and Overcooked tasks.
@@ 2 Related Work
Training AHT Agents. The training of ego agent policies that near-optimally collaborate with diverse previously unseen teammates has been explored in AHT [50]. Most AHT methods follow the two-stage design process, where the generation of a fixed training set of teammate policies is followed by AHT training. Given teammates from the training set, AHT methods [34] train an ego agent to model teammates [3] by first identifying their important characteristics (e.g., goals, beliefs, policies) based on their observed behavior, and then estimating the best-response policy to these teammates based on the inferred characteristics. Recent AHT methods [40, 39, 64, 55] use neural networks trained using reinforcement learning [48, 35]. To further improve AHT training, several approaches learn a distribution for sampling teammate policies during training based on maximizing the worst-case returns [54] or regret [19, 12] of trained agents. While few, exceptions to the two-stage process include methods that expect to encounter a continual stream of teammates at deployment [37, 38, 62], and self-play based methods, which do not explicitly optimize for diversity [60, 14].
Teammate Generation for AHT & ZSC. Recent work removes the need to predefine teammate policy sets by generating diverse teammates during or before agent training. Other-Play [22] creates symmetry-equivalent teammates while training the agent policy, while E3T [60] mixes the agent’s current policy with a random policy to encourage diversity. FCP [51] trains teammates via repeated CMARL runs with different seeds, later improved by methods maximizing information-theoretic diversity objectives such as Jensen-Shannon divergence [33], mutual information [32], and entropy [59, 63]. More recent approaches [10, 42, 62] generate teammates that require distinct best-response strategies by maximizing adversarial diversity metrics, similar to ROTATE’s cooperative regret. Unlike ROTATE, these methods (i) maximize regret between generated teammates rather than with the trained agent, (ii) fix the teammate set prior to training, and (iii) evaluate regret only at the initial state. This last property leads to sabotaging teammates that harm cooperation in states unseen in self-play, motivating heuristic solutions in prior work [16, 46], and a systematic objective in ROTATE.
Open-Ended Learning (OEL). OEL [28, 52] studies algorithms that continually generate novel tasks to train generally capable agents [23, 4]. Many OEL approaches in RL take the form of unsupervised environment design (UED) [17], which improves generalization by designing or sampling new environments with varied initial states. Some methods directly train neural networks to propose environments that induce high regret in the agent [17], while others selectively sample curated tasks generated by procedural generators based on criteria such as expected return [56], TD-error [25], regret [24], or learnability [44]. In competitive MARL, OEL often produces new opponents through self-play [49, 29]. For AHT, MACOP [62] generates novel teammates via an adversarial diversity objective optimized with evolutionary methods and similar to the objectives studied by by Charakorn et al. [10] and  Rahman et al. [41]. Thus, the objective can yield sabotaging teammates when applied only to the initial state. In contrast, ROTATE adopts a more systematic training objective that we demonstrate leads to performance gains.
@@ 3 Background
The interaction between agents in an AHT setting may be modeled as a decentralized Markov decision process (Dec-MDP) [5]. A Dec-MDP is characterized by a 7-tuple, $\langle N,S,\{\mathcal{A}^{i}\}_{i=1}^{|N|},P,p_{0},R,\gamma\rangle$, where $N$, $S$, and $\gamma$ respectively denote the index set of agents within an interaction, the state space, and a discount rate, $0\leq\gamma\leq 1$. Every interaction between agents begins from a state sampled from the initial state distribution, $s_{0}\sim p_{0}(s)$. At timestep $t$, each agent, $i\in{N}$, jointly executes an action selected from its action space, $a^{i}_{t}\in\mathcal{A}^{i}$, based on the observed state, $s_{t}$, and its policy, $\pi^{i}(s^{i}_{t})$. We assume that teammates choose their actions only based on the current state. Meanwhile, the AHT agent, also referred to as the ego agent, selects actions based on its state-action history, which is necessary to distinguish between different types of teammates effectively. Denoting the set of all probability distributions over a set $X$ as $\Delta(X)$, the execution of the joint action, $a_{t}=(a^{1}_{t},\dots,a_{t}^{|N|})$, results in agents observing a new state, $s_{t+1}$, sampled according to the environment transition function, $P:S\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{|N|}\mapsto\Delta{(S)}$, and a common scalar reward, $r_{t}$, based on the reward function, $R:S\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{|N|}\mapsto\mathbb{R}$.
@@ 4 Ad Hoc Teamwork Problem Formulation
AHT methods aim to train an adaptive policy that an ego agent can follow to achieve maximal return when collaborating with an unknown set of evaluation teammates. Formalizing the interaction between agents as a Dec-MDP, this section outlines the objective of AHT. While the most general AHT setting considers a possibly varying number of ego agents and teammates within an interaction [55, 40], this formalization addresses the more straightforward case where there is only a single ego agent within a team.
Let $\pi^{\text{ego}}$ refer to the ego agent’s policy, and $\pi^{-i}$ denote the |N|1|N|-1 policies of the AHT agent’s teammates. We denote the returns of an ego agent that follows $\pi^{\text{ego}}$ to collaborate with teammates controlled by $\pi^{-i}$, starting from state $s$, as:
$V(s|\pi^{-i},\pi^{\text{ego}})=\mathbb{E}_{\begin{subarray}{c}a^{\text{ego}}_{t}\sim\pi^{\text{ego}},\\ a^{-i}_{t}\sim\pi^{-i},P\end{subarray}}\Bigg[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\Bigg|s_{0}=s\Bigg].$ (1)
@@
$\text{BR}(\pi^{-i})\in\underset{\pi}{\operatorname*{arg\,max}\ }\mathbb{E}_{s\sim p_{0}}\left[V(s|\pi,\pi^{-i})\right].$ (3)
In some cases, AHT algorithms can estimate this optimal policy by using $\Pi^{\text{eval}}$ to train an ego agent policy that maximizes $V(\psi^{\text{eval}},\Pi^{\text{eval}},\pi^{\text{ego}})$ when $\Pi^{\text{eval}}$ is known. [^1] In the context of reinforcement-learning-based AHT algorithms, “known” means that an AHT algorithm has unlimited sampling access to the teammate policies. However, most AHT methods address the more challenging case where $\Pi^{\text{eval}}$ is unknown, which is the setting that this paper adopts as well. While our methods assume no knowledge of $\Pi^{\text{eval}}$ during training, we follow standard practice [39, 40, 64, 55] by manually designing a diverse $\Pi^{\text{eval}}$ for evaluation purposes, as we later describe in Section 7.
When $\Pi^{\text{eval}}$ is unknown, AHT algorithms [34] learn by interacting with policies from the training set, $\Pi^{\text{train}}$, which are learned or manually designed by leveraging an expert’s domain knowledge about the characteristics of $\Pi^{\text{eval}}$. After forming the set of training teammates, current AHT algorithms use RL to discover an ego agent policy based on interactions with joint policies sampled from $\Pi^{\text{train}}$. While the precise training objective varies with the AHT algorithm, methods commonly estimate the ego agent policy maximizing the expected return during interactions with joint policies sampled uniformly from $\Pi^{\text{train}}$, which we describe below:
@@
Naturally, even $\pi^{*,\text{ego}}(\Pi^{\text{train}})$ may be suboptimal with respect to $\Pi^{\text{eval}}$ and $\psi^{\text{eval}}$, due to the potential distribution shift caused by differences between the training and evaluation objectives.